OpenStudy (anonymous):

What is the value of the expression 1/4 ^-3 ? A. 12 B. 64 C. 1/64 D. 1/12

OpenStudy (anonymous):

@Compassionate @tHe_FiZiCx99

OpenStudy (smileyxl3):

It's 64

OpenStudy (superdavesuper):

\[x ^{-y}=(1/x)^{y}\]

OpenStudy (anonymous):

will u help me with 2 more

OpenStudy (anonymous):

What is the value of the expression 2/4^-3 ?

OpenStudy (smileyxl3):

128

OpenStudy (anonymous):

What value of k solves the equation? k^-3 = 1/27

OpenStudy (whpalmer4):

\[k^{-3} = \frac{1}{27}\]\[k^{-3} = \frac{1}{k^3}\]\[\frac{1}{k^3} = \frac{1}{27}\]any ideas about what \(k\) is?

OpenStudy (anonymous):

ummm -81

OpenStudy (anonymous):

-9

OpenStudy (whpalmer4):

well, for \[\frac{1}{k^3} = \frac{1}{27}\]we need to have \[k^3 = k*k*k= 27\]right?

OpenStudy (anonymous):

yea

OpenStudy (whpalmer4):

can you think of a number that makes that true? 1*1*1 = 1 2*2*2 = 8

OpenStudy (anonymous):

its 3

OpenStudy (whpalmer4):

bingo!

OpenStudy (anonymous):

or -3???

OpenStudy (whpalmer4):

good question, but no, not if there is an even power involved. \[x^2=9\] for both \(x=3\) and \(x=-3\) however, we have \[x^3=27\] \[3*3*3 = 27\] \[(-3)*(-3)*(-3) = -27\]

OpenStudy (whpalmer4):

sorry, "not if there is an ODD power involved"

OpenStudy (anonymous):

What value of n solves the equation 4^n + 1/64?

OpenStudy (whpalmer4):

that's not an equation, as there is no equals sign :-)

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

What value of n solves the equation 4^n = 1/64

OpenStudy (whpalmer4):

\[4^n = \frac{1}{64}\] Any ideas on whether \(n\) is positive or negative?

OpenStudy (anonymous):

positive

OpenStudy (whpalmer4):

well, sadly, no :-) if it positive, then we are going to have something like \[4*4*4 = \frac{1}{64}\] (for some number of 4's) But that's not ever going to be true, is it?

OpenStudy (whpalmer4):

No, we need a negative exponent, because we need it to look something like \[\frac{1}{4}*\frac{1}{4}*... = \frac{1}{64}\]right?

OpenStudy (anonymous):

yea?

OpenStudy (whpalmer4):

A negative exponent would let us do this: \[4^n = \frac{1}{4^m} = \frac{1}{64}\](if \(m = -n\))

OpenStudy (whpalmer4):

So that makes our problem be \[4^m = 64\]How many times do you have to multiply 4 with itself to get 64? 4=4 4*4=16 4*4*4=64 4*4*4*4=256 etc.

OpenStudy (anonymous):

okay thanks will u help me with some more???

OpenStudy (whpalmer4):

So what is the answer for that problem? \[4^n = \frac{1}{64}\]\[n=\]